# Simplicial arrangements with few double points

### v1

@@ 1. Introduction
Grünbaum’s catalogue of combinatorial equivalence classes contains three infinite families $\mathcal{R}(0)$, $\mathcal{R}(1)$ and $\mathcal{R}(2)$, see below. It also contains about one hundred sporadic arrangements (the most complicated of them is formed by $37$ lines). A comprehensible description of the sporadic arrangements is given in [5]. The list is not completely stable since Cuntz recently discovered a new sporadic arrangement (see [2]).
@@
Note that arrangements of families $\mathcal{R}(1)$ and $\mathcal{R}(2)$ are projectively rigid. In other words, two combinatorially equivalent arrangement are identical up to a projective transformation. For completeness we prove this fact in Corollary 4.4, deducing it from an elementary and more general Proposition 4.3.
@@
Suppose that $\mathcal{A}$ is an arrangement of $n$ lines in $\mathbb{RP}^{2}$. Let $K>0$ be a real parameter. Suppose that at most $Kn$ vertices of $\mathcal{A}$ are double points and that $n\geq exp\leavevmode\nobreak\ exp(cK^{c})$ for some constant $c$ independent from $K$. Then, up to a projective transformation, $\mathcal{A}$ differs by at most $O(K)$ lines (which can be added or deleted) from an example of one of the following three types:
@@
A geometric interpretation of the few double points hypothesis is as follows. For each simplicial arrangement, we can construct a polyhedral metric on $\mathbb{RP}^{2}$ by gluing equilateral triangles according to the incidence structure of the arrangement. Double points will be vertices of positive curvature. Since the total curvature is a topological invariant (Gauss-Bonnet equality), the number of double points controls the number of vertices of order at least four (vertices of negative curvature).
If there is a small number of double points, then most vertices in the arrangements are triple points and large domains of the arrangement will be covered by an equilateral lattice. This argument has been developed by Melchior in his proof of Sylvester-Gallai theorem, see [6].
It appears that case (1) of Theorem 1.6 is realized by the arrangements of family $\mathcal{R}(0)$ while case (2) is realized by the regular simplicial arrangements of families $\mathcal{R}(1)$ and $\mathcal{R}(2)$. In Section 3.3, we prove that no curve that is projectively dual to an irreducible cubic can be tangent to most lines of a simplicial arrangement, eliminating case (3) of Theorem 1.6 for simplicial arrangements.
@@
Finally, we prove in Section 4 that a simplicial arrangement cannot differ from a regular arrangement by only few lines unless it is actually a regular arrangement (up to the addition of the line at infinity in the case of arrangements obtained from a regular polygon with an even number of sides). This is the third and final step in the proof of Theorem 1.5.
@@ 2.1. Conventions
The next result follows from Gauss-Bonnet formula applied to the polyhedral metric on $\mathbb{RP}^{2}$ obtained by replacing every face of a simplicial arrangement by an equilateral triangle.
Lemma 2.2.
For any simplicial arrangement, we have $\sum\limits_{k\geq 2}v_{k}(k-3)=-3$ where $v_{k}$ the number of vertices incident to $k$ lines of the arrangement.
Lemma 2.2 implies in particular that every simplicial arrangement contains at least three double points. It also motivates the convention according to which vertices of higher order are the vertices of order at least four.
@@ 2.2. Exceptional properties of family $\mathcal{R}(0)$
Near-pencils, or equivalently simplicial arrangements of combinatorial classes $\mathcal{A}(k+1,0)$ for $k\geq 2$ are very special among simplicial arrangements.
@@
Assuming $A$ is a vertex of order $k$, $2k$ edges of arrangement $\mathcal{A}$ are incident to $A$. Two of them are incident to $B$ while the $22k-$ 2 other edges are incident to double points. Unless $A$ is a double point, two double points incident to $A$ are adjacent. Since the arrangement is simplicial, Lemma 2.3 implies that the arrangement belongs to family $\mathcal{R}(0)$. The same reasoning applies to $B$. If every vertex of $\mathcal{A}$ is a double point, then Lemma 2.2 implies that the only vertices of the arrangement are three double points and we obtain arrangement $\mathcal{A}(3,0)$. ∎
@@ 2.3. Star of a vertex
Since there is no pair of adjacent double points in an arrangement that does not belong to $\mathcal{R}(0)$ (see Lemma 2.3), every line contains at most two exterior edges of $S(C)$. The number $d$ of double points among exterior vertices of $S(C)$ thus satisfies $d\leq k$. Since a line contains one or two exterior edges of $S(C)$ depending whether the common exterior vertex is a double point or not, subarrangement $\mathcal{F}_{C}$ contains at least $2k-d$ distinct lines. ∎
A basic result on geometry of lines cutting out a convex polygon will be useful for our further classification works.
@@
If an arrangement $\mathcal{A}$ of $m\geq 3$ lines cuts out a convex $m$-gon $\mathcal{P}$ in $\mathbb{RP}^{2}$, then it is combinatorially equivalent to the arrangement formed by $m$ tangent lines of a unit circle in such a way that the tangency points form a regular $m$-gon. In particular, every vertex of $\mathcal{A}$ is a double point and the faces are polygon $\mathcal{P}$, $m$ triangles adjacent to $\mathcal{P}$ and $\frac{m(m-3)}{2}$ quadrilaterals.
@@ 3. Estimates on lines tangent to a dual cubic in a simplicial arrangement
The main goal of this section is to prove Proposition 1.7. Throughout the section we denote by $\hat{\Gamma}$ a cubic curve in $\mathbb{RP}^{2}$ and by $\Gamma$ the projectively dual curve. By definition, $\Gamma$ is formed by all lines tangent to $\hat{\Gamma}$, where for a singular point $x\in\hat{\Gamma}$ all lines containing $x$ are tangent to $\hat{\Gamma}$. Taking multiplicities into account, $\Gamma$ is defined by a degree $6$ polynomial. We start by recalling a qualitative description of irreducible cubics and their duals.
3.1. Description of dual cubics
Recall that every irreducible planar cubic curve $\hat{\Gamma}$ can be written in Weierstrass form $y^{2}=x^{3}+ax+b$. It can have at most one singularity, namely, a double point, a cusp or an isolated point (see Theorem 8.4 in [1]). This defines three classes of planar curves projectively dual to a singular irreducible cubic:
@@ 3.1. Description of dual cubics
Figure 2. Nodal and acnodal cubics and their duals.
@@
Examples of irreducible cubics and their dual curves are given in Figures 2 and 3.
Figure 3. Smooth irreducible cubics and their duals.
@@ 3.2. Families of lines decomposing a convex polygon into triangles
We define again $\mathcal{G}^{\prime}$ as the set of lines of $\mathcal{G}$ that intersect the interior of $\mathcal{P}^{\prime}$ and =$\mathcal{F}^{\prime}=\{{\Lambda_{2},\dots,\Lambda_{|\mathcal{F}|}\}}$. Then $(\mathcal{P}^{\prime},\mathcal{F}^{\prime},\mathcal{G}^{\prime})$ is a counter-example such that $|\mathcal{F}^{\prime}|=|\mathcal{F}|-1$ and $|\mathcal{G}^{\prime}|\leq|\mathcal{G}|-1$. This contradicts minimality of $(\mathcal{P},\mathcal{F},\mathcal{G})$. ∎
@@
We stress that in Proposition 1.7 and below a line of $\mathcal{A}$ is considered to be tangent to $\Gamma$ only if its dual point belongs to $\hat{\Gamma}$. We will split the proof of the proposition into five cases depending on the shape of the irreducible cubic $\hat{\Gamma}$ (see Section 3.1).
@@
We follow exactly the same reasoning as in the case where $\hat{\Gamma}$ is smooth and connected. This time $\Gamma$ is the union of the cuspidal cubic $y^{2}=x^{3}$ with the line at infinity. The curvilinear triangle $\Delta$ is formed by the union of the cuspidal cubic and the segment $x\geq 0$ on the line at infinity. The only difference with the first case is that the straight side of $\Delta$ is tangent to at most one line of $\mathcal{A}$ the one containing it. ∎
@@
Adding (3.1) to (3.2) multiplied by $6$, we get $217g\geq k_{\Delta}+k_{O}-21=k-$ 21. Using $g=n-k$, we finally get $n\geq\frac{8}{7}k-3$. ∎
@@
4. Rigidity of regular simplicial arrangements and proof of the main theorem
@@ 4.1. Geometry of regular simplicial arrangements
A regular simplicial arrangement is the union of two subarrangements:
  • $\mathcal{A}_{\gamma}$ formed by the sides of a regular $m$-gon (all the lines are tangent to a circle $\gamma$);
  • a pencil $\mathcal{A}_{C}$ formed by $m$ lines containing the centre of circle $\gamma$.
In view of our future needs, we give a precise geometric description of the line arrangement $\mathcal{A}_{\gamma}$. We assume from now on that $\mathcal{A}_{\gamma}$ lies on the complex plane in such a way that circle $\gamma$ is the unit circle and lines $L_{1},\dots,L_{m}$ have their tangency point with $\gamma$ at the $m^{th}$ roots of the unit $e^{\frac{2ij\pi}{m}}$ where $j\in\mathbb{Z}/m\mathbb{Z}$.
@@
Suppose $A,B,C$ belong to the same line and none of them is at infinity. Then considering the triangle $AOC$, we get the following expression for the length of bisector $OB$
@@
In our case, we have $\theta=\frac{\pi}{m}$, $OA=\frac{1}{\cos((k-3)\theta)}$, $OB=\frac{1}{\cos(k\theta)}$ and $OC=\frac{1}{\cos((k+3)\theta)}$ where $k=|j-l|$. We obtain
@@
In the remaining case, $\Lambda$ intersects the radial sectors of $N_{3}$ (it follows from the previous arguments that $\Lambda$ also intersects the radial sectors of $N_{1}$ and $N_{5}$). We check that if a radial sector of $N_{3}$ intersected by $\Lambda$ belongs to a triangle of arrangement $\mathcal{A}_{\gamma}$, then $N_{2}$ (or $N_{4}$ depending on the orientation we chose) belongs to the edge of this triangle that is incident to polygon $\mathcal{P}$ and $N_{1}$ (or $N_{5}$) cannot be a vertex of $\mathcal{A}_{\gamma}$. The same problem appears if the radial sector of $N_{3}$ containing $\Lambda$ belongs to polygon $\mathcal{P}$. Hence the radial sectors of $N_{3}$ intersected by $\Lambda$ are contained in quadrilaterals of arrangement $\mathcal{A}_{\gamma}$. The arguments given previously show similarly that the four edges $[N_{1},N_{2}]$, $[N_{2},N_{3}]$, $[N_{3},N_{4}]$ and $[N_{4},N_{5}]$ are contained in quadrilaterals of $\mathcal{A}_{\gamma}$.
@@
where $|j-l|\geq 4$ (otherwise, $[N_{1},N_{5}]$ would intersect the interior of a triangle of $\mathcal{A}_{r}$ or the interior of $\mathcal{P}$). Lemma 4.1 rules out these cases. We conclude that there is no such configuration $N_{1},N_{2},N_{3},N_{4},N_{5}$. ∎
4.2. Projective rigidity of regular simplicial arrangements

### v2

@@ 1. Introduction
Grünbaum’s catalogue of combinatorial equivalence classes contains three infinite families $\mathcal{R}(0)$, $\mathcal{R}(1)$ and $\mathcal{R}(2)$, see below. It also contains about one hundred sporadic arrangements (the most complicated of them is formed by $37$ lines). A comprehensible description of the sporadic arrangements is given in [8]. Thanks to the work of Cuntz, we now know that this list is not entirely complete. A computer verification of simplicial arrangements with at most 27 lines revealed four previously unknown examples (see [3]). More recently, a new sporadic arrangement consisting of 35 lines was discovered (see [4]).
@@
Note that arrangements of families $\mathcal{R}(1)$ and $\mathcal{R}(2)$ are projectively rigid. In other words, two combinatorially equivalent arrangement are identical up to a projective transformation. This has been proved in Theorem 3.6 and Corollary 3.7 of [2]. For completeness we prove this fact in the Appendix in Corollary A.2, deducing it from an elementary and more general Proposition A.1.
@@
Suppose that $\mathcal{A}$ is an arrangement of $n$ lines in $\mathbb{RP}^{2}$. Let $K>0$ be a real parameter. Suppose that at most $Kn$ vertices of $\mathcal{A}$ are double points and that $n\geq exp\penalty 10000\ exp(cK^{c})$ for some constant $c$ independent from $K$. Then, up to a projective transformation, $\mathcal{A}$ differs by at most $O(K)$ lines (which can be added or deleted) from an example of one of the following three types:
@@
This theorem says that arrangements with a linear bound on the number of double points are governed by cubic curves. Namely we see that in all three cases (1), (2), (3) when we take the arrangement of points dual to the lines of $\mathcal{A}$, most of such points lie on some cubic curve. The case (1) is a bit degenerate, and dual points lie on a line, in case (2) they lie on a disjoint union of a conic and a line, and in case (3) on an irreducible cubic curve. As for the arrangement $\mathcal{A}$ itself, most of its lines are tangent to the curve projectively dual to a cubic curve (again cases (1) and (2) are a bit degenerate here).
Case (1) of Theorem 1.6 is realized by the arrangements of family $\mathcal{R}(0)$ while case (2) is realized by the regular simplicial arrangements of families $\mathcal{R}(1)$ and $\mathcal{R}(2)$. However in Section 4.2, we prove that no curve that is projectively dual to an irreducible cubic can be tangent to most lines of a simplicial arrangement, eliminating case (3) of Theorem 1.6 for simplicial arrangements.
@@
Finally, we prove in Section 5 that a simplicial arrangement cannot differ from a regular arrangement by only few lines unless it is actually a regular arrangement (up to the addition of the line at infinity in the case of arrangements obtained from a regular polygon with an even number of sides). This is the final step in the proof of Theorem 1.5.
@@
Data availability statement. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
@@ 2.2. Exceptional properties of family $\mathcal{R}(0)$
Near-pencils, or equivalently simplicial arrangements of combinatorial classes $\mathcal{A}(k+1,0)$ for $k\geq 2$ are very special among simplicial arrangements. The following result appears as Lemma 3.2 in [3] in the more general context of simplicial wirings, where it is presented without proof as it is considered folklore. We include a proof here for the sake of completeness.
@@
Assuming $A$ is a vertex of order $k$, $2k$ edges of arrangement $\mathcal{A}$ are incident to $A$. Two of them are incident to $B$ while the 2k22k-2 other edges are incident to double points. Unless $A$ is a double point, two double points incident to $A$ are adjacent. Since the arrangement is simplicial, Lemma 2.2 implies that the arrangement belongs to family $\mathcal{R}(0)$. The same reasoning applies to $B$. If every vertex of $\mathcal{A}$ is a double point, then Lemma 2.2 again proves that the arrangement belongs to $\mathcal{R}(0)$. In other words, it is a near-pencil where every line except one is incident to a point (which is a double point). In this case, we obtain arrangement $\mathcal{A}(3,0)$. ∎
@@ 2.3. Star of a vertex
Since there is no pair of adjacent double points in an arrangement that does not belong to $\mathcal{R}(0)$ (see Lemma 2.2), every line contains at most two exterior edges of $S(C)$. The number $d$ of double points among exterior vertices of $S(C)$ thus satisfies $d\leq k$. Since a line contains one or two exterior edges of $S(C)$ depending whether the common exterior vertex is a double point or not, subarrangement $\mathcal{F}_{C}$ contains at least 2kd2k-d distinct lines. ∎
Each arrangement cuts $\mathbb{RP}^{2}$ into several convex polygons. It turns out that in case the number of sides of one of these polygons coincides with the number of lines of the arrangement, the arrangement has a particularly simple form.
@@
If one face of an arrangement $\mathcal{A}$ of $m\geq 3$ lines is a convex $m$-gon $\mathcal{P}$ in $\mathbb{RP}^{2}$, then $\mathcal{A}$ is combinatorially equivalent to the arrangement formed by $m$ tangent lines of a unit circle in such a way that the tangency points form a regular $m$-gon. In particular, every vertex of $\mathcal{A}$ is a double point and the faces are polygon $\mathcal{P}$, $m$ triangles adjacent to $\mathcal{P}$ and $\frac{m(m-3)}{2}$ quadrilaterals.
@@
3. Recollection on cubic curves and their duals
In this short section, we give a qualitative description of real cubic curves and their duals. We denote by $\hat{\Gamma}$ a cubic curve in $\mathbb{RP}^{2}$ and by $\Gamma$ the projectively dual curve. By definition, $\Gamma$ is formed by all lines tangent to $\hat{\Gamma}$, where for a singular point $x\in\hat{\Gamma}$ all lines containing $x$ are tangent to $\hat{\Gamma}$. Taking multiplicities into account, $\Gamma$ is defined by a degree $6$ polynomial.
Recall that every irreducible planar cubic curve $\hat{\Gamma}$ can be written in Weierstrass form $y^{2}=x^{3}+ax+b$. An irreducible cubic curve can have at most one singularity, namely, a double point, a cusp or an isolated point (see Theorem 8.4 in [1]), this happens when the polynomial $x^{3}+ax+b$ has a double root. Altogether, we have $3$ such irreducible cubic curves up to an isomorphism:
@@ 3. Recollection on cubic curves and their duals
Nodal Acnodal Smooth Smooth
connected disconnected
Primary [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Dual [Uncaptioned image] [Uncaptioned image]    [Uncaptioned image] [Uncaptioned image]
Table 1. Some cubic curves and their projective duals
Examples of irreducible cubic curves and their dual curves are given in Table 1. Here the cubic curve $\hat{\Gamma}$ is called primary, and the dual cubic curve $\Gamma$ is called dual. We don’t provide the figure for the easiest case - the semi-cubic parabola.
The main goal of this section is to prove Proposition 1.7.
@@ 4.1. Families of lines decomposing a convex polygon into triangles
We define again $\mathcal{G}^{\prime}$ as the set of lines of $\mathcal{G}$ that intersect the interior of $\mathcal{P}^{\prime}$ and $\mathcal{F}^{\prime}=\{{\Lambda_{2},\dots,\Lambda_{|\mathcal{F}|}\}}$. Then $(\mathcal{P}^{\prime},\mathcal{F}^{\prime},\mathcal{G}^{\prime})$ is a counter-example such that $|\mathcal{F}^{\prime}|=|\mathcal{F}|-1$ and $|\mathcal{G}^{\prime}|\leq|\mathcal{G}|-1$. This contradicts minimality of $(\mathcal{P},\mathcal{F},\mathcal{G})$. ∎
@@
We stress that in Proposition 1.7 and below a line of $\mathcal{A}$ is considered to be tangent to $\Gamma$ only if its dual point belongs to $\hat{\Gamma}$. We will split the proof of the proposition into five cases depending on the shape of the irreducible cubic $\hat{\Gamma}$ (see Section 3).
@@
We follow exactly the same reasoning as in the case where $\hat{\Gamma}$ is smooth and connected. This time $\Gamma$ is the union of the cuspidal cubic $y^{2}=x^{3}$ with the line at infinity. The curvilinear triangle $\Delta$ is formed by the union of the cuspidal cubic and the segment $x\geq 0$ on the line at infinity. The only difference with the first case is that the straight side of $\Delta$ is tangent to at most one line of $\mathcal{A}$ - the one containing it. ∎
@@
Adding (4.1) to (4.2) multiplied by $6$, we get $7g\geq k_{\Delta}+k_{O}-21=k-21$. Using g=nkg=n-k, we finally get $n\geq\frac{8}{7}k-3$. ∎
@@
5. Proof of the main theorem
@@ 5.1. Geometry of regular simplicial arrangements
A regular simplicial arrangement is the union of two subarrangements defined with respect to a circle $\gamma$ which is, for our purpose, the unit circle of the complex plane:
  • $\mathcal{A}_{\gamma}$ formed by the lines $L_{1},\dots,L_{m}$ that are tangent to $\gamma$ at the $m^{th}$ roots of the unit $e^{\frac{2ij\pi}{m}}$ where $j\in\mathbb{Z}/m\mathbb{Z}$;
  • a pencil $\mathcal{A}_{C}$ formed by the $m$ lines containing the centre of circle $\gamma$ and a $2m^{th}$ root of the unit $e^{\frac{ij\pi}{m}}$ where $j\in\mathbb{Z}/m\mathbb{Z}$;.
@@
Suppose $A,B,C$ belong to the same line and none of them is at infinity. Then considering the triangle $AOC$, we get the following expression for the length of bisector OBOB
@@
In our case, we have $\theta=\frac{\pi}{m}$, $OA=\frac{1}{\cos((k-3)\theta)}$, $OB=\frac{1}{\cos(k\theta)}$ and $OC=\frac{1}{\cos((k+3)\theta)}$ where k=|jl|k=|j-l|. We obtain
@@
In the remaining case, $\Lambda$ intersects the radial sectors of $N_{3}$ (it follows from the previous arguments that $\Lambda$ also intersects the radial sectors of $N_{1}$ and $N_{5}$). We check that if a radial sector of $N_{3}$ intersected by $\Lambda$ belongs to a triangle of arrangement $\mathcal{A}_{\gamma}$, then $N_{2}$ (or $N_{4}$ depending on the orientation we chose) belongs to the edge of this triangle that is incident to polygon $\mathcal{P}$ and $N_{1}$ (or $N_{5}$) cannot be a vertex of $\mathcal{A}_{\gamma}$. The same problem appears if the radial sector of $N_{3}$ containing $\Lambda$ belongs to polygon $\mathcal{P}$. Hence the radial sectors of $N_{3}$ intersected by $\Lambda$ are contained in quadrilaterals of arrangement $\mathcal{A}_{\gamma}$. The arguments given previously show similarly that the four edges $N_{1},N_{2}]$, $N_{2},N_{3}]$, $N_{3},N_{4}]$ and $N_{4},N_{5}]$ are contained in quadrilaterals of $\mathcal{A}_{\gamma}$.
@@
where $|j-l|\geq 4$ (otherwise, $N_{1},N_{5}]$ would intersect the interior of a triangle of $\mathcal{A}_{r}$ or the interior of $\mathcal{P}$). Lemma 5.1 rules out these cases. We conclude that there is no such configuration $N_{1},N_{2},N_{3},N_{4},N_{5}$. ∎
Proposition 5.3.
Consider a simplicial arrangement $\mathcal{A}$ of $n\geq 15$ lines in $\mathbb{RP}^{2}$ that differs by at most $k<\frac{1}{13}n-\frac{14}{13}$ lines from a regular simplicial arrangement $\mathcal{A}_{r}$ formed by the sides and symmetry axes of a regular $m$-gon $\mathcal{P}$ ($\mathcal{A}_{r}$ belongs to the combinatorial class $\mathcal{A}(2m,1)$). Then $\mathcal{A}$ coincides either with $\mathcal{A}_{r}$ or with the union of $\mathcal{A}_{r}$ with the line at infinity.
Lemma 5.4.
Consider a simplicial arrangement $\mathcal{A}$ of $n\geq 15$ lines in $\mathbb{RP}^{2}$ that differs by at most $k<\frac{1}{7}n+\frac{1}{7}$ lines from a regular simplicial arrangement $\mathcal{A}_{r}$ formed by the sides and symmetry axes of a regular $m$-gon $\mathcal{P}$ ($\mathcal{A}_{r}$ belongs to the combinatorial class $\mathcal{A}(2m,1)$). Then no line $\Lambda$ of $\mathcal{A}\setminus\mathcal{A}_{r}$ contains the centre $C$ of the regular polygon $\mathcal{P}$. Besides, $\mathcal{A}$ cannot be a near-pencil.
Proof.
By our assumptions, $\frac{n-k}{2}\leq m\leq\frac{n+k}{2}$. Also arrangement $\mathcal{A}$ contains at least mkm-k sides of polygon $\mathcal{P}$. Since $m-k\geq\frac{n-3k}{2}>\frac{2}{7}n-\frac{3}{14}\geq 4$ (because $k<\frac{1}{7}n+\frac{1}{7}$ and $n\geq 15$), there is a polygon with at least four sides formed by the union of several faces of $\mathcal{A}$. Hence it cannot belong to family $\mathcal{R}(0)$.
Assume by contradiction there exists a line $\Lambda\in\mathcal{A}\setminus\mathcal{A}_{r}$ containing the centre $C$ of the regular polygon $\mathcal{P}$. Let $M_{1},\dots,M_{t}$ be a cyclically ordered set of points on $\Lambda$ where $M_{1}=C$ while $M_{2},\dots,M_{t}$ are the intersection points (other than $C$) of $\Lambda$ with the other lines of the symmetric difference $\mathcal{A}\Delta\mathcal{A}_{r}$. We have $t\leq k$ since $|\mathcal{A}\Delta\mathcal{A}_{r}|\leq k$ and $\Lambda\in\mathcal{A}\Delta\mathcal{A}_{r}$. A line containing $C$ cannot contain any other vertex of $\mathcal{A}_{r}$ without being itself a line of $\mathcal{A}_{r}$. Therefore, $\Lambda$ crosses at most one line of $\mathcal{A}_{r}$ at each of points of $M_{2},\dots,M_{t}$. To sum up, $\Lambda$ intersects $m$ lines of $\mathcal{A}_{r}$ in $C=M_{1}$, at most t1t-1 lines of $\mathcal{A}_{r}$ in $M_{2},\dots,M_{t}$ and k1k-1 lines of $\mathcal{A}\Delta\mathcal{A}_{r}$ in $M_{2},\dots,M_{t}$. It follows that $\Lambda$ intersects at most m+t+k2m+t+k-2 lines of $\mathcal{A}$ in $M_{1},\dots,M_{t}$.
By hypothesis, inside any open interval $M_{i},M_{i+1}[$, line $\Lambda$ only intersects lines of $\mathcal{A}_{r}$. Besides, it cannot contain any vertex of $\mathcal{A}_{r}$ (because $C=M_{1}$ and $\Lambda$ cannot contain any other vertex of $\mathcal{A}_{r}$ without being a line of $\mathcal{A}_{r}$). It follows that every vertex of $M_{i},M_{i+1}[$ is a double point of $\mathcal{A}$. We know from Lemma 2.2 that two double points cannot be adjacent in a simplicial arrangement that does not belong to family $\mathcal{R}(0)$. It follows that in any open interval $M_{i},M_{i+1}[$, $\Lambda$ intersects at most one line of $\mathcal{A}$.
We conclude that line $\Lambda$ intersects at most m+t+k2m+t+k-2 lines of $\mathcal{A}$ in $M_{1},\dots,M_{t}$ and at most $t$ lines of $\mathcal{A}$ outside $M_{1},\dots,M_{t}$. Thus, we have $n-1\leq m+2t+k-2\leq\frac{1}{2}n+\frac{7}{2}k-2$ (because $t\leq k$ and $m\leq\frac{n+k}{2}$). It follows that $\frac{1}{7}n+\frac{1}{7}\leq k$. This contradicts our assumptions, and so there is no $\Lambda\in\mathcal{A}\setminus\mathcal{A}_{r}$ containing $C$. ∎
Lemma 5.5.
Proof.
  • a)
  • b)
Let’s show first that $t\leq k+1$. Indeed, line $\Lambda$ has at most k1k-1 intersection points with the lines of $\mathcal{A}\Delta\mathcal{A}_{r}$, so there are at most k1k-1 points of type a) among $M_{i}$’s. And at most $2$ of $M_{i}$’s are of type b), since $\mathcal{A}_{r}$ has exactly $2m$ double points, namely the midpoints of edges of the convex polygon $\mathcal{P}$.
Next, we claim that there are at most $4t$ lines of ${\mathcal{A}}\setminus\Lambda$ that are incident to $M_{1},\ldots,M_{t}$. Indeed, since $C\notin\Lambda$, each $M_{i}$ is at most a triple point of $\mathcal{A}_{r}$, and $|\mathcal{A}\Delta\mathcal{A}_{r}\setminus\Lambda|=k-1$ . Taking the sum of contributions, and using $t\leq k+1$ we get $3t+k-1\leq 4t$.
Our goal will now be to prove that each open interval $M_{i},M_{i+1}[$ of $\Lambda$ intersects at most $9$ lines of $\mathcal{A}\setminus\Lambda$. If the set $\{{M_{1},\dots,M_{t}\}}$ is empty, we will prove similarly that the whole projective line intersects at most $9$ lines of $\mathcal{A}\setminus\Lambda$. Then, counting the lines of $\mathcal{A}\setminus\Lambda$ intersected by $\Lambda$ inside and outside of $M_{1},\dots,M_{t}$, we obtain that $n-1\leq 4t+9t\leq 13k+13$ (since $t\leq k+1$). This contradicts the inequality $k<\frac{1}{13}n-\frac{14}{13}$ and $\Lambda$ is therefore the line at infinity.
It remains to prove that in each open interval $M_{i},M_{i+1}[$ (or the whole line if $\{{M_{1},\dots,M_{t}\}}$ is empty), $\Lambda$ intersects at most $9$ lines of $\mathcal{A}\setminus\{{\Lambda\}}$. We assume by contradiction that there is such an open interval (that can be the whole line) $M_{i},M_{i+1}[$ in $\Lambda$ that intersects at least $10$ lines of $\mathcal{A}$. Since the intersections of $\Lambda$ with the lines of $\mathcal{A}\Delta\mathcal{A}_{r}$ are contained in $M_{1},\dots,M_{n}$, all these lines of $\mathcal{A}$ that $\Lambda$ intersects in $M_{i},M_{i+1}[$ actually are lines of $\mathcal{A}_{r}$. It also follows from the definition of $M_{1},\dots,M_{t}$ that every line of $\mathcal{A}_{r}$ intersecting $\Lambda$ outside of $M_{1},\dots,M_{t}$ actually belongs to $\mathcal{A}$.
The only vertices of $\mathcal{A}_{r}$ that line $\Lambda$ can meet outside of $M_{1},\dots,M_{t}$ are triple points of $\mathcal{A}_{r}$. Therefore, the vertices of $\mathcal{A}$ contained in the open segment $M_{i},M_{i+1}[$ are either quadruple points of $\mathcal{A}$ (if they are vertices of $\mathcal{A}_{r}$) or double points of $\mathcal{A}$ (otherwise).
Arrangement $\mathcal{A}$ does not belong to family $\mathcal{R}(0)$ (see the last claim in Lemma 5.4). Hence, there cannot be two consecutive double points in $M_{i},M_{i+1}[$ (see Lemma 2.2). We prove now that there cannot be two consecutive quadruple points $A,B$ of $\mathcal{A}$ in $M_{i},M_{i+1}[$. Indeed, such segment $A,B]\subset]M_{i},M_{i+1}[$ is drawn between two vertices of $\mathcal{A}_{r}$ so unless it is contained in a line of $\mathcal{A}_{r}$ (and $\Lambda\notin\mathcal{A}_{r}$), $S$ intersects the interior of an edge of $\mathcal{A}_{r}$. This is a contradiction with the statement that $A,B$ are consecutive intersection points with $\mathcal{A}_{r}$ in $M_{i},M_{i+1}[$. So vertices of segment $M_{i},M_{i+1}[$ alternate between double and quadruple points. Since $M_{i},M_{i+1}[$ intersects at least $10$ lines of $\mathcal{A}_{r}$, we conclude that the open segment $M_{i},M_{i+1}[$ contains a sequence of five consecutive vertices $N_{1},N_{2},N_{3},N_{4},N_{5}$ where $N_{1},N_{3},N_{5}$ are quadruple points while $N_{2},N_{4}$ are double points.
Since $N_{1},N_{3},N_{5}$ are triple points in $\mathcal{A}_{r}$ while $N_{2},N_{4}$ are interior points of edges in $\mathcal{A}_{r}$, Proposition 5.2 proves that there is no line containing this configuration of five points. This final contradiction ends the proof. ∎
Proof of Proposition 5.3.
  • (i)
  • (ii)
  • (iii)
    $\frac{m-3}{2}$ if $m$ is odd.
We conclude that among quadrilaterals crossed by $\Lambda$, at least $\frac{m-5}{2}$ have the interior disjoint from the line at infinity. Since $\mathcal{A}$ is simplicial, each of these quadrilaterals is contained in a triangle of $\mathcal{A}$, so at least one of their sides should also be removed from $\mathcal{A}_{max}$. Since the same line can contain a side of at most two of these quadrilaterals, the total number of lines of $\mathcal{A}_{\max}\setminus\mathcal{A}$ is at least $\frac{m-5}{4}$. We obtain $k\geq\frac{m-5}{4}$. Since we have $n\leq 2m+1$, we get $k\geq\frac{n}{8}-\frac{11}{8}$. Since $n\geq 15$, this contradicts $k<\frac{1}{13}n-\frac{14}{13}$.
We conclude that the the center of polygon $\mathcal{P}$ is incident to exactly $m$ lines of $\mathcal{A}$. It follows from Proposition 2.5 that the star $S(C)$ of $C$ in the simplicial arrangement $\mathcal{A}$ is cut out by at least $m$ lines (Lemma 5.4 proves that $\mathcal{A}$ cannot belong to family $\mathcal{R}(0)$). We deduce that $\mathcal{A}$ coincides with $\mathcal{A}_{max}$. ∎
5.3. Proof of the main theorem
Combining the results from Sections 2.3, 4 and 5, we finally prove the asymptotic classification of simplicial arrangements under the additional hypothesis of a linear bound on the number of double points.
Proof of Theorem 1.5.
Assuming that a simplicial arrangement $\mathcal{A}$ of $n$ lines contains at most $Kn$ double points (for some $K>1$) and that $n\geq exp\penalty 10000\ exp(cK^{c})$, Theorem 1.6 proves that $\mathcal{A}$ differs (up to a projective transformation) by at most $O(K)$ lines from one of the three models described in the statement of the theorem.
In the first case, $\mathcal{A}$ differs by a small proportion of lines from a pencil. In particular, provided constant $c$ is large enough, $\mathcal{A}$ contains a vertex $C$ of order $k_{C}$ arbitrarily close to $n$. Proposition 2.5 proves that if $\mathcal{A}$ is not a simplicial arrangement of family $\mathcal{R}(0)$, the star $S(C)$ of $C$ is cut out by at least $k_{C}$ lines that do not contain $C$ and $k_{C}\leq\frac{n}{2}$. It follows that $\mathcal{A}$ belongs to family $\mathcal{R}(0)$.
Appendix A Projective rigidity of regular simplicial arrangements